OpenStudy (anonymous):

Generate the first 5 terms of this sequence: f(1) = 0 and f(2) = 1, f(n) = f(n - 1) + f(n - 2), for n > 2. 0, -1, 1, 0, 2 0, 1, 1, 2, 3 0, 1, 2, 2, 3 0, 1, 1, 2, 2

OpenStudy (anonymous):

@Hero

OpenStudy (mertsj):

f(3)=f(3-1)+f(3-2) f(3)=f(2)+f(1) f(3)=1+0=1

OpenStudy (anonymous):

what is the answer @Mertsj

OpenStudy (mertsj):

You need to find the f(4) and the f(5) then you will have the first 5 terms because terms 1 and 2 are given. I did term 3 for you. You need to do the next 2.

OpenStudy (anonymous):

i do not know how to do it @Mertsj

OpenStudy (mertsj):

Do you see that your problem says : f(n)=f(n-1)+f(n-2) ?????

OpenStudy (anonymous):

No I do not may you just help me with the other two problems, because I am not going to get anywhere right now @Mertsj